Let F, G be meromorphic functions such that Fn+Gn=1, assuming n≥4 if necessary. Prove that F and G are constants.


1
Expert's answer
2021-07-12T17:17:15-0400

Let us first consider a function "h:z\mapsto (F(z),G(z))" this function takes its values in the complex curve "\mathcal{C}\subseteq \bar\mathbb{C}^2" defined by an equation "z^4+w^4=1". This curve is a compact complex curve of genus "\frac{1}{2} (n-1)(n-2)" which is "\geq 2" for "n\geq 4". By the Riemann uniformization theorem, a compact complex curves of genus "\geq 2" admit a universal covering by a unit disc "D". Therefore, there exists a lifting "\bar{h}: \mathbb{C}\to D" and by the Liouville's theorem it is constant (as it is a bounded entire function). As "h=\pi\circ \bar h", where "\pi" is the covering map of the curve "\mathcal{C}", we have "h=const" and thus "F(z),G(z)=const".

The condition "n\geq 4" is not only sufficient, but necessary - for "n=1" there are obvious linear solutions, for "n=2" we can take trigonometric functions and for "n=3" there are non-trivial solutions using the elliptic functions.